# Eigenvalue analysis of three-state quantum walks with general coin matrices

### v4

@@ Abstract
Mathematical analysis of the spectral properties of the time evolution operator in quantum walks is essential for understanding key dynamical behaviors such as localization and long-term evolution. The inhomogeneous three-state case, in particular, poses substantial analytical challenges due to its higher internal degrees of freedom and the absence of translational invariance. We develop a general framework for the spectral analysis of three-state quantum walks on the one-dimensional lattice with arbitrary time evolution operators. Our novel approach is based on a transfer matrix formulation that reduces the infinite-dimensional eigenvalue problem to a tractable system of two-dimensional recursions, enabling exact characterization of localized eigenstates. This framework applies broadly to space-inhomogeneous models, including those with finite defects and two-phase structures. We rigorously derive necessary and sufficient conditions for the existence of discrete eigenvalues, along with a complete description of the corresponding eigenvalues and eigenstates, which are known to underlie quantum localization phenomena. Using this method, we perform exact numerical analyses of the Fourier walk with spatial inhomogeneity, revealing the emergence of localization despite its delocalized nature in the homogeneous case. Our results provide mathematical tools and physical insights into the structure of quantum walks, offering a systematic path for identifying and characterizing localized quantum states in complex quantum systems.
@@ I Introduction
A vital aspect of studying QWs is analyzing the spectral properties of the time evolution operator. This analysis provides insights into the behavior and features of the walk, thereby enhancing our understanding of the underlying dynamics [19, 20, 18]. A distinctive feature of infinite-dimensional systems, unlike their finite-dimensional counterparts, is the coexistence of continuous spectra and discrete eigenvalues, which reflect different physical behaviors. This structure allows for rigorous mathematical analysis [14, 20]. Discrete eigenvalues in infinite systems typically correspond to localized states and are key to understanding topological phenomena in QWs. In topological phases, discrete eigenvalues within spectral gaps often indicate the presence of topologically protected bound states, which arise due to the bulk-edge correspondence [1, 2]. Unlike extended bulk states, these bound states remain robust under system variations as long as the topological invariant remains unchanged. Additionally, QWs can exhibit dynamically induced topological transitions, where spectral gaps close and reopen, signaling a shift in the topological phase. Thus, the study of discrete eigenvalues in infinite QWs provides deep insights into topologically protected phenomena and phase transitions, revealing features akin to those in topological insulators and superconductors. A similar discussion arises in the context of quantum algorithm speedup [7, 8, 9, 21]. In quantum algorithm design, the key objective is to determine how quickly the initial state evolves into the desired state. QWs used for algorithms are typically defined on finite-dimensional systems, and the speedup depends on the distinct eigenvalues and the spectral gap separating them from the rest of the spectrum [22, 23, 24, 25]. A higher overlap between the initial and desired states with an eigenstate of a distinct eigenvalue, along with a large spectral gap, often enables faster transitions. This is crucial for achieving quantum speedup and is analogous to how strongly and quickly the QW localizes to the desired state.
@@
Thus, in this paper, we focus on a mathematically explicit approach to analyze the (discrete) eigenvalues and eigenvectors of the time evolution operator. This rigorous analysis provides deeper insights into the spectral structure of the system and facilitates understanding of phenomena such as localization and the long-term behavior of QWs. It is well-known that the occurrence of localization is equivalent to the existence of eigenvalues of the time evolution operator. Moreover, a quantitative investigation of localization can be conducted by deriving the time-averaged limit distribution using the eigenvalues and eigenvectors [18]. The transfer matrix method stands as a powerful tool for eigenvalue analysis and has been applied to several types of QW models by author [17, 26], and even for non-unitary QWs [27].
@@
In this article, we consider the eigenvalue analysis on the space-inhomogeneous three-state QWs on a one-dimensional integer lattice. Unlike the standard two-state QW, which has been extensively analyzed, the three-state QW exhibits fundamentally different spectral properties, leading to richer and more complex dynamics [28, 29]. This distinction makes three-state QWs particularly intriguing and has led to intensive study in various articles across different contexts [16, 28, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. We also specifically consider the two-phase model with a finite number of defects, which encompasses important and typical space-inhomogeneous models such as the one-defect model and the two-phase model [44, 17]. These models have been studied extensively as they serve as crucial examples in understanding the impact of inhomogeneity on QWs and provide a rich landscape for investigating localization phenomena. Previous studies [26] have conducted eigenvalue analysis for three-state QWs with a restricted class of coin matrices, including Grover matrices. In contrast, our work significantly broadens the scope by establishing a general framework that applies to arbitrary unitary coin matrices, demonstrating that precise eigenvalue characterization is achievable well beyond previously studied cases.
@@
As the main theorem, we establish a necessary and sufficient condition for the eigenvalue equation of the time evolution operator in three-state QWs with general coin matrices, providing an explicit characterization of both the eigenvalues and the corresponding eigenvectors. However, calculating explicit eigenvalues analytically for a general coin matrix remains challenging due to the complexity of solving the associated algebraic equations. Therefore, we rely on numerical calculations to demonstrate the practicality of our method. For our numerical analysis, we focus on one-defect and two-phase Fourier walks. The Fourier walk has been extensively studied in the context of space-homogeneous models [45, 46, 47], where the coin matrices are identical at all positions. Notably, unlike the Grover walk, it has been proven that the Fourier walk does not exhibit localization in the homogeneous case [47]. However, our numerical results show that the Fourier walk exhibits localization in both the one-defect and two-phase models.
@@
The remainder of this paper is organized as follows. In Section II, we define our three-state QWs with a self-loop on the integer lattice, incorporating both the coin and shift operators. Additionally, we explicitly define the concept of localization. Section III introduces the transfer matrix framework for a general coin matrix and establishes methods for addressing the eigenvalue problem. We demonstrate how the eigenvalue problem can be expressed using transfer matrices, and Theorem 1 provides a necessary and sufficient condition for solving the eigenvalue problem of two-phase QWs with finite defects. Section IV focuses on the numerical analysis of the Fourier walk. We include visual representations of the numerical results for the eigenvalue equation and the probability distribution, illustrating the occurrence of localization and the existence of eigenvalues.
II Definition
@@
Let $\{C_{x}\}_{x\in\mathbb{Z}\cup\{\pm\infty\}}$ be a sequence of $3\times 3$ unitary matrices, which is given by:
$C_{x}=\left[\begin{array}[]{ c c c }a_{x}^{(1,1)}&a_{x}^{(1,2)}&a_{x}^{(1,3)}\\ a_{x}^{(2,1)}&a_{x}^{(2,2)}&a_{x}^{(2,3)}\\ a_{x}^{(3,1)}&a_{x}^{(3,2)}&a_{x}^{(3,3)}\end{array}\right].$
@@
The QW is governed by a time evolution operator $U$, and understanding its eigenvalues is essential for exploring phenomena such as localization, where the quantum walker tends to remain confined to a particular region of the lattice. The degree of localization of the walker is highly dependent on the initial state, and it is quantified by the limit distribution, which relies heavily on the overlap between the initial state and the eigenstates of $U$ [18]. In this article, however, we focus on whether the model is capable of exhibiting localization, which implies the existence of an eigenstate that could overlap with the initial state.
@@
III Eigenvalue Analysis with the Transfer Matrix
@@ III.1 Reducing the dimensionality
III.2 Eigenvalue analysis via transfer matrix
@@
In the next Proposition 1, we give the main statement for this subsection, where we convert the eigenvalue equation to a reduced transfer matrix equation. Note that, when $A_{x}(\lambda)=0$, which also means $a_{x}^{(1,1)}e^{i\lambda}=e^{i\Delta_{x}}\overline{a_{x}^{(3,3)}}$, we cannot construct a transfer matrix. In this case, we must treat it separately. The following proposition rewrites the statement of Lemma 1 using the transfer matrix. The proof of this proposition is provided in Appendix A, there we show that $A_{x}(\lambda)=0$ implies $D_{x}(\lambda)=0$, thus the Eqs. 5 and 6 gives the direct relation $\tilde{\Psi}_{1}(x)$ to $\tilde{\Psi}_{2}(x)$ and $\tilde{\Psi}_{1}(x+1)$ to $\tilde{\Psi}_{2}(x+1)$ for $A_{x}(\lambda)=0$ case, as detailed below.
@@
$\bullet$ For $x$ with $A_{x}(\lambda)=0$:
$a_{x}^{(3,3)}\overline{a_{x}^{(2,3)}}\tilde{\Psi}_{1}(x+1)=\overline{a_{x}^{(1,1)}}a_{x}^{(1,2)}\tilde{\Psi}_{2}(x+1).$ (7)
@@
III.3 Eigenvalue analysis for two-phase QW with defects
@@
$C_{x}=\begin{cases}C_{\infty},&x\in[x_{+},\infty),\\ C_{x},&x\in[x_{-},x_{+}),\\ C_{-\infty},&x\in(-\infty,x_{-}),\end{cases}$
where $x_{+}\geq 0,\ x_{-}<0$. We say the model with this condition as two-phase QW with finite number of defects. As a remark, we can easily replace $C_{\pm\infty}$ with periodic coin matrices by following the previous research [48]. However, we opt not to explore this generalization as it leads to notably more complex notation.
@@
For all $\lambda\in[-\pi,\pi)$ and $x\in\mathbb{Z}\cup\{\pm\infty\}$, we have
@@
In this paper, we define the square root of the complex number $a=|a|e^{i\theta},\ \theta\in[0,2\pi)$ as $\sqrt{a}=\sqrt{|a|}e^{i\frac{\theta}{2}}$. Consequently, we designate $\zeta_{x}^{<}$ as one of $\zeta_{x}^{\pm}$ such that $\left|\zeta_{x}^{<}\right|\leq 1$, and $\zeta_{x}^{>}$ such that $\left|\zeta_{x}^{>}\right|\geq 1.$ Additionally, we introduce |$\ket{v_{x}^{>}}$ and |$\ket{v_{x}^{<}}$ as normalized eigenvector of $T_{x}$ corresponding to $\zeta_{x}^{>}$ and $\zeta_{x}^{<}$, respectively. For $\tilde{\Psi}(x)$ to decay as $x\to\pm\infty$ it is necessary that both $|\zeta_{\pm\infty}^{>}|$ and $|\zeta_{\pm\infty}^{<}|$ are not equal to $1$. We define the set $\Lambda$ as the set of $\lambda$ for which this condition is satisfied.
@@
For $\lambda\in[0,2\pi)$, $e^{i\lambda}$ is an eigenvalue of $U$ if and only if there exists $\tilde{\Psi}:\mathbb{Z}\rightarrow\mathbb{C}^{2}$ such that Conditions 1, 2, and 3 are all satisfied.
@@
$\circ$ Case 2. $A_{x}(\lambda)=0$: There exists $k_{1},k_{2}\in\mathbb{C}$ such that
$\tilde{\Psi}(x)=k_{1}\begin{bmatrix}a_{x}^{(3,3)}\overline{a_{x}^{(3,2)}}\\ \overline{a_{x}^{(1,1)}}a_{x}^{(2,1)}\end{bmatrix},\ \tilde{\Psi}(x+1)=k_{2}\begin{bmatrix}\overline{a_{x}^{(1,1)}}a_{x}^{(1,2)}\\ a_{x}^{(3,3)}\overline{a_{x}^{(2,3)}}\end{bmatrix}.$
@@
$\circ$ Case 1. $A_{\infty}(\lambda)\neq 0$: $|\mathrm{tr}(T_{\infty}(\lambda))|>2$ and
$T_{\infty}(\lambda)\tilde{\Psi}(x_{+})=\zeta_{\infty}^{<}\tilde{\Psi}(x_{+}).$
$\tilde{\Psi}(x_{+})=k\begin{bmatrix}a_{\infty}^{(3,3)}\overline{a_{\infty}^{(3,2)}}\\ \overline{a_{\infty}^{(1,1)}}a_{\infty}^{(2,1)}\end{bmatrix},$
and for $x>x_{+},$ there exists $k_{x}\in\mathbb{C}$ such that
@@
$\circ$ Case 1. $A_{-\infty}(\lambda)\neq 0$: $|\mathrm{tr}(T_{-\infty}(\lambda))|>2$ and
$T_{-\infty}(\lambda)\tilde{\Psi}(x_{-})=\zeta_{-\infty}^{>}\tilde{\Psi}(x_{-}).$
$\ \tilde{\Psi}(x_{-})=k\begin{bmatrix}\overline{a_{-\infty}^{(1,1)}}a_{-\infty}^{(1,2)}\\ a_{-\infty}^{(3,3)}\overline{a_{-\infty}^{(2,3)}}\end{bmatrix},$
and for $x<x_{-},$ there exists $k_{x}\in\mathbb{C}$, such that
$\ \tilde{\Psi}(x)=\begin{cases}k_{x}\begin{bmatrix}a_{-\infty}^{(3,3)}\overline{a_{-\infty}^{(3,2)}}\\ \overline{a_{-\infty}^{(1,1)}}a_{-\infty}^{(2,1)}\end{bmatrix}&\left(\frac{a_{-\infty}^{(3,3)}}{\overline{a_{-\infty}^{(1,1)}}}\right)^{2}=\frac{a_{-\infty}^{(1,2)}a_{-\infty}^{(2,1)}}{\overline{a_{-\infty}^{(3,2)}}\overline{a_{-\infty}^{(2,3)}}},\\ \mathbf{0}&otherwise.\end{cases}$
@@
$\Lambda_{0}=\left\{e^{i\lambda}\in\mathbb{C}\mid\exists x\in\mathbb{Z},\ A_{x}(\lambda)=0\right\}.$

### v5

# Spectral analysis of three-state quantum walks with general coin matrices

@@ Abstract
Mathematical analysis of the spectral properties of the time evolution operator in quantum walks is essential for understanding key dynamical behaviors such as localization and long-term evolution. The inhomogeneous three-state case, in particular, poses substantial analytical challenges due to its higher internal degrees of freedom and the absence of translational invariance. We develop a general framework for the spectral analysis of three-state quantum walks on the one-dimensional lattice with arbitrary time evolution operators. Our approach is based on a transfer matrix formulation that reduces the infinite-dimensional eigenvalue problem to a tractable system of two-dimensional recursions, enabling exact characterization of eigenstates. This framework applies broadly to space-inhomogeneous models, including those with finite defects and two-phase structures. We rigorously derive necessary and sufficient conditions for the existence of point spectrum, along with a complete description of the corresponding eigenvalues and eigenstates, which are known to underlie quantum localization phenomena. Furthermore, we give a complete spectral decomposition—discrete spectrum, flat-band eigenvalues (of infinite multiplicity), and absolutely continuous spectrum—with explicit characterization of each component. Using this method, we perform exact numerical analyses of the Fourier walk with spatial inhomogeneity, revealing the emergence of localization despite its delocalized nature in the homogeneous case. Our results provide mathematical tools and physical insights into the structure of quantum walks, offering a systematic path for identifying and characterizing localized quantum states in complex quantum systems.
@@ I Introduction
A vital aspect of studying QWs is analyzing the spectral properties of the time evolution operator. This analysis provides insights into the behavior and features of the walk, thereby enhancing our understanding of the underlying dynamics [19, 20, 18]. A distinctive feature of infinite-dimensional systems, unlike their finite-dimensional counterparts, is the coexistence of continuous spectra and discrete eigenvalues, which reflect different physical behaviors. This structure allows for rigorous mathematical analysis [14, 20]. Discrete eigenvalues in infinite systems typically correspond to localized states and are key to understanding topological phenomena in QWs. In topological phases, discrete eigenvalues within spectral gaps often indicate the presence of topologically protected bound states, which arise due to the bulk-edge correspondence [1, 2]. Unlike extended bulk states, these bound states remain robust under system variations as long as the topological invariant remains unchanged. Additionally, QWs can exhibit dynamically induced topological transitions, where spectral gaps close and reopen, signaling a shift in the topological phase. Thus, the study of discrete eigenvalues and continuous spectra in infinite QWs provides deep insights into topologically protected phenomena and phase transitions, revealing features akin to those in topological insulators and superconductors. A similar discussion arises in the context of quantum algorithm speedup [7, 8, 9, 21]. In quantum algorithm design, the key objective is to determine how quickly the initial state evolves into the desired state. QWs used for algorithms are typically defined on finite-dimensional systems, and the speedup depends on the distinct eigenvalues and the spectral gap separating them from the rest of the spectrum [22, 23, 24, 25]. A higher overlap between the initial and desired states with an eigenstate of a distinct eigenvalue, along with a large spectral gap, often enables faster transitions. This is crucial for achieving quantum speedup and is analogous to how strongly and quickly the QW localizes to the desired state.
@@
Thus, in this paper, we focus on a mathematically explicit approach to analyze the (discrete) eigenvalues and eigenvectors of the time evolution operator. This rigorous analysis provides deeper insights into the spectral structure of the system and facilitates understanding of phenomena such as localization and the long-term behavior of QWs. It is well-known that the occurrence of localization is equivalent to the existence of eigenvalues of the time evolution operator. Moreover, a quantitative investigation of localization can be conducted by deriving the time-averaged limit distribution using the eigenvalues and eigenvectors [18]. The transfer matrix method stands as a powerful tool for eigenvalue analysis and has been applied to several types of QW models by author [17, 26], and even for non-unitary QWs [27]. Beyond identifying discrete eigenvalues, this paper also reveals the structure of the essential spectrum. This complete the spectral picture which is crucial for understanding the interplay between localization, delocalization, and spectral degeneracy. In particular, resolving the essential spectrum allows us to determine the presence or absence of spectral gaps, which govern the robustness of localized modes. This level of spectral resolution makes our analysis substantially more comprehensive than previous studies.
@@
In this article, we consider the spectral analysis on the space-inhomogeneous three-state QWs on a one-dimensional integer lattice. Unlike the standard two-state QW, which has been extensively analyzed, the three-state QW exhibits fundamentally different spectral properties, leading to richer and more complex dynamics [28, 29]. This distinction makes three-state QWs particularly intriguing and has led to intensive study in various articles across different contexts [16, 28, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. We also specifically consider the two-phase model with a finite number of defects, which encompasses important and typical space-inhomogeneous models such as the one-defect model and the two-phase model [44, 17]. These models have been studied extensively as they serve as crucial examples in understanding the impact of inhomogeneity on QWs and provide a rich landscape for investigating localization phenomena. Previous studies [26] have conducted eigenvalue analysis for three-state QWs with a restricted class of coin matrices, including Grover matrices. In contrast, our work significantly broadens the scope by establishing a general framework that applies to arbitrary unitary coin matrices, demonstrating that precise eigenvalue characterization is achievable well beyond previously studied cases. In contrast to homogeneous two-state walks, the three-state model generically exhibits flat-band eigenvalues of infinite multiplicity, which form part of the essential spectrum and correspond to compactly supported modes originating from algebraic symmetries of the tail coins. Our transfer-matrix framework makes this structure explicit: by analyzing the asymptotic transfer matrices, we determine the absolutely continuous part of the spectrum and distinguish it from flat-band (infinite multiplicity) contributions and defect-induced (finite multiplicity) point spectra.
@@
As the main theorem, we establish a necessary and sufficient condition for the eigenvalue equation of the time evolution operator in three-state QWs with general coin matrices, providing an explicit characterization of both the eigenvalues and the corresponding eigenvectors. In addition, we characterize the essential spectrum to obtain a full description of the spectral structure. However, calculating explicit eigenvalues analytically for a general coin matrix remains challenging due to the complexity of solving the associated algebraic equations. Therefore, we rely on numerical calculations to demonstrate the practicality of our method. For our numerical analysis, we focus on one-defect and two-phase Fourier walks. The Fourier walk has been extensively studied in the context of space-homogeneous models [45, 46, 47], where the coin matrices are identical at all positions. Notably, unlike the Grover walk, it has been proven that the Fourier walk does not exhibit localization in the homogeneous case [47]. However, our numerical results show that the Fourier walk exhibits localization in both the one-defect and two-phase models.
@@
The remainder of this paper is organized as follows. In Section II, we define our three-state QWs with a self-loop on the integer lattice, incorporating both the coin and shift operators. After we explicitly define the concept of localization, we introduce the transfer matrix framework for a general coin matrix and establishes methods for addressing the eigenvalue problem. In Section III, we state our main results for the spectral analysis on two-phase QWs with finite defects. Theorem 1 provides a necessary and sufficient condition for solving the eigenvalue problem. Furthermore, Theorem 2 establishes the complete spectral decomposition, providing a full classification of discrete, flat-band eigenvalue with infinite multiplicity, and absolutely continuous components of the spectrum. Section IV focuses on the numerical analysis of the Fourier walk. We include visual representations of the numerical results for the eigenvalue equation and the probability distribution, illustrating the occurrence of localization and the existence of eigenvalues.
II Three state quantum walk and transfer matrix
II.1 Definition of three-state quantum walk
@@
Let $\{\mathsf{C}_{x}\}_{x\in\mathbb{Z}\cup\{\pm\infty\}}$ be a sequence of $3\times 3$ unitary matrices, which is given by:
$\mathsf{C}_{x}=\left[\begin{array}[]{ c c c }a_{x}^{(1,1)}&a_{x}^{(1,2)}&a_{x}^{(1,3)}\\ a_{x}^{(2,1)}&a_{x}^{(2,2)}&a_{x}^{(2,3)}\\ a_{x}^{(3,1)}&a_{x}^{(3,2)}&a_{x}^{(3,3)}\end{array}\right].$
@@
The QW is governed by a time evolution operator $U$, and understanding its eigenvalues is essential for exploring phenomena such as localization, where the quantum walker tends to remain confined to a particular region of the lattice. The degree of localization of the walker is highly dependent on the initial state, and it is quantified by the limit distribution, which relies heavily on the overlap between the initial state and the eigenstates of $U$ [18].
@@
II.2 Eigenvalue analysis
@@ II.2.1 Reducing the dimensionality
II.2.2 Transfer matrix on reduced space
@@
$T_{x}(\lambda)=\frac{1}{a_{x}^{(1,1)}e^{i\lambda}-e^{i\Delta_{x}}\overline{a_{x}^{(3,3)}}}\begin{bmatrix}e^{i\lambda}\left(e^{i\lambda}-a_{x}^{(2,2)}\right)&-a_{x}^{(1,3)}e^{i\lambda}-e^{i\Delta_{x}}\overline{a_{x}^{(3,1)}}\\ a_{x}^{(3,1)}e^{i\lambda}+e^{i\Delta_{x}}\overline{a_{x}^{(1,3)}}&-e^{i\Delta_{x}}\left(e^{-i\lambda}-\overline{a_{x}^{(2,2)}}\right)\end{bmatrix}$. (7)
Here, $e^{i\Delta_{x}}$ is a determinant of the coin matrix $\det\mathsf{C}_{x}$. In the next Proposition 1, we give the main statement for this subsection, where we convert the eigenvalue equation to a reduced transfer matrix equation. Note that, when $A_{x}(\lambda)=0$, which also means $a_{x}^{(1,1)}e^{i\lambda}=e^{i\Delta_{x}}\overline{a_{x}^{(3,3)}}$, we cannot construct a transfer matrix. In this case, we must treat it separately. The following proposition rewrites the statement of Lemma 1 using the transfer matrix. The proof of this proposition is provided in Appendix A, there we show that $A_{x}(\lambda)=0$ implies $D_{x}(\lambda)=0$, thus the Eqs. 5 and 6 gives the direct relation $\tilde{\Psi}_{1}(x)$ to $\tilde{\Psi}_{2}(x)$ and $\tilde{\Psi}_{1}(x+1)$ to $\tilde{\Psi}_{2}(x+1)$ for $A_{x}(\lambda)=0$ case, as detailed below.
@@
$\bullet$ For $x$ with $A_{x}(\lambda)=0$: there exists $k_{x},k^{\prime}_{x}\in\mathbb{C}$ such that
$\tilde{\Psi}(x)=k_{x}v_{x},\quad\tilde{\Psi}(x+1)=k^{\prime}_{x}w_{x},$
where
$v_{x}:=\begin{bmatrix}a_{x}^{(3,3)}\overline{a_{x}^{(3,2)}}\\ \overline{a_{x}^{(1,1)}}a_{x}^{(2,1)}\end{bmatrix},\quad w_{x}:=\begin{bmatrix}\overline{a_{x}^{(1,1)}}a_{x}^{(1,2)}\\ a_{x}^{(3,3)}\overline{a_{x}^{(2,3)}}\end{bmatrix}$.
@@
III Spectral analysis for two-phase QW with defects
@@
$\mathsf{C}_{x}=\begin{cases}\mathsf{C}_{\infty},&x\in[x_{+},\infty),\\ \mathsf{C}_{x},&x\in[x_{-},x_{+}),\\ \mathsf{C}_{-\infty},&x\in(-\infty,x_{-}),\end{cases}$
where $x_{+}\geq 0,\ x_{-}<0$. This implies that the coin operator is spatially homogeneous in the regions $x\geq x_{+}$ and $x<x_{-}$, while allowing for arbitrary coin matrices in the finite interval between $x_{-}$ and $x_{+}$. We say the model with this condition as two-phase QW with finite number of defects. As a remark, we can easily replace $C_{\pm\infty}$ with periodic coin matrices by following the previous research [48]. However, we opt not to explore this generalization as it leads to notably more complex notation.
III.1 Eigenvalue and Eigenvector
@@
For all $\lambda\in[-\pi,\pi)$ and $x\in\mathbb{Z}\cup\{\pm\infty\}$, the transfer matrix given by (7) satisfies
@@
[…]
@@
For $\lambda\in[0,2\pi)$, $e^{i\lambda}$ is an eigenvalue of $U$ if and only if there exists non-zero $\tilde{\Psi}:\mathbb{Z}\rightarrow\mathbb{C}^{2}$ such that Conditions 1, 2, and 3 are all satisfied.
@@
$\circ$ Case 2. $A_{x}(\lambda)=0$: There exists $k_{x},k^{\prime}_{x}\in\mathbb{C}$ such that
$\tilde{\Psi}(x)=k_{x}v_{x},\ \tilde{\Psi}(x+1)=k^{\prime}_{x}w_{x}.$
@@
$\circ$ Case 1. $A_{\infty}(\lambda)\neq 0$: $\tilde{\Psi}(x_{+})=\mathbf{0}$ or
$|\mathrm{tr}(T_{\infty}(\lambda))|>2,\ T_{\infty}(\lambda)\tilde{\Psi}(x_{+})=\zeta_{\infty}^{<}\tilde{\Psi}(x_{+}).$
$\circ$ Case 2. $A_{\infty}(\lambda)=0$: There exists $k_{x_{+}}\in\mathbb{C}$ such that $\tilde{\Psi}(x_{+})=k_{x_{+}}v_{\infty},$ and for $x>x_{+},$ there exists $k_{x}\in\mathbb{C}$ such that
$\tilde{\Psi}(x)=\begin{cases}k_{x}v_{\infty}&\left(\frac{a_{\infty}^{(3,3)}}{\overline{a_{\infty}^{(1,1)}}}\right)^{2}=\frac{a_{\infty}^{(1,2)}a_{\infty}^{(2,1)}}{\overline{a_{\infty}^{(3,2)}}\overline{a_{\infty}^{(2,3)}}},\\ \mathbf{0}&otherwise.\end{cases}$
@@
$\circ$ Case 1. $A_{-\infty}(\lambda)\neq 0$: $\tilde{\Psi}(x_{-})=\mathbf{0}$ or
$|\mathrm{tr}(T_{-\infty}(\lambda))|>2,\ T_{-\infty}(\lambda)\tilde{\Psi}(x_{-})=\zeta_{-\infty}^{>}\tilde{\Psi}(x_{-}).$
$\circ$ Case 2. $A_{-\infty}(\lambda)=0$: There exists $k^{\prime}_{x_{-}-1}\in\mathbb{C},$ such that $\tilde{\Psi}(x_{-})=k^{\prime}_{x_{-}-1}w_{-\infty},$ and for $x<x_{-},$ there exists $k_{x}\in\mathbb{C}$, such that
[…]
@@
Furthermore, we have the following result on the multiplicity of eigenvalues.
Proposition 2.
The multiplicity of $e^{i\lambda}$ as an eigenvalue of $U$ is described as follows.
(i) $\dim\ker(U-e^{i\lambda})=\infty$ if and only if one of the following holds:
$A_{\infty}(\lambda)=0,\ \left(\frac{a^{\infty}_{33}}{a^{\infty}_{11}}\right)^{2}=\frac{a^{\infty}_{12}a^{\infty}_{21}}{a^{\infty}_{32}a^{\infty}_{23}}$
$A_{-\infty}(\lambda)=0,\ \left(\frac{a^{-\infty}_{33}}{a^{-\infty}_{11}}\right)^{2}=\frac{a^{-\infty}_{12}a^{-\infty}_{21}}{a^{-\infty}_{32}a^{-\infty}_{23}}.$
(ii) If the condition of (i) does not hold, then the multiplicity is finite and satisfies
$\dim\ker(U-e^{i\lambda})\ \leq\ 1+N(\lambda),$
where $N(\lambda)$ is the number of exceptional sites in the interval $x_{-},x_{+})$:
$N(\lambda):=|\{x\in[x_{-},x_{+})\;:\;A_{x}(\lambda)=0\}|.$
Equality holds exactly when, for every exceptional site $x$, there exists an eigenstate for which the coefficients $k_{x}$ and $k^{\prime}_{x}$ in Condition 1, Case 2 of Theorem 1 are both nonzero.
Proof.
By Proposition 1, if $A_{x}(\lambda)\neq 0$ then $\tilde{\Psi}$ is uniquely propagated by $\tilde{\Psi}(x+1)=T_{x}(\lambda)\tilde{\Psi}(x)$ and no new freedom arises. If $A_{x}(\lambda)=0$, the second case replaces this by $\tilde{\Psi}(x)=k_{x}u_{x}$ and $\tilde{\Psi}(x+1)=k^{\prime}_{x}w_{x}$, so each exceptional site can contribute (at most) one extra degree of freedom, which is realized exactly when an eigenvector exists with $k_{x}\neq 0$ and $k^{\prime}_{x}\neq 0$.
If an exceptional site occurs in a tail ($A_{\pm\infty}(\lambda)=0$) and the corresponding algebraic condition in (i) holds, then Theorem 1 yields infinitely many linearly independent eigenvectors from infinitely many choices of $k_{x}\in\mathbb{C}$, hence $\dim\ker(U-e^{i\lambda})=\infty$. Otherwise, the only exceptional sites contribute to the multiplicity lie in the finite defect region $x_{-},x_{+})$. Each of the $N(\lambda)$ exceptional sites can contribute at most one additional independent choice, when an eigenvector exists with $k_{x}\neq 0$ and $k^{\prime}_{x}\neq 0$, giving $\dim\ker(U-e^{i\lambda})\leq 1+N(\lambda)$. ∎
@@
$\Lambda_{0}=\left\{e^{i\lambda}\in\mathbb{C}\mid\exists x\in\mathbb{Z},\ A_{x}(\lambda)=0\right\}.$ (8)